Depreciation 200

Exports 150

public transfer 200

Gross domestic private investment 300

corporate income tax 100

factor receipts from abroad 400

gevenment expenditure on goods and services 250

interest income 800

compensation of employees 2600

net interest on government debit 50

indirect business taxes 100

factor payments to abroad 200

imports 200

proprietors incom(profits) 700

retained corporate profit 200

personal consumption expenditure 4500

personal taxes 100

social security contribution 50

rental income 600. From the above data find personale disposable income.


1
Expert's answer
2021-09-02T10:48:00-0400

Personal Disposable Income is the difference between personal income and personal income tax.

 "Personal\space Disposable\space Income =Personal\space Income − Personal\space Income\space Tax"


From the given information, 

Personal Income includes "=Rental \space Income + Proprietor's\space Income + Social\space Security \space Contribution + Interest \space Income + Compensation\space of \space employees"

Personal Income "= 600+700+50+800+2600"

Personal Income "= 4750"


Personal Income tax includes personal taxes and indirect business tax.

Personal Income Tax "=100+100"

Personal Income Tax "= 200"

 

Thus,Personal Disposable Income "= 4750-200"

Personal Disposable Income"= 4450"

 

 

Answer: Personal Disposable Income "=\$4450."